Complete the following table to see how decimals and negative numbers work. X + Y XY 2,420.7 3,841.4 2,700.0 -3,100.0
Answer #1

X=2420.7, Y= 3841.1, X+Y = 2420.7 + 3841.4= 6262.1

X=2700.0, Y= -3100, X+Y = 2700.0 + (-3100.0)= -400.0